OpenStudy (anonymous):

need steps to solve this...please help! 10 * log(4x)=25

OpenStudy (anonymous):

log 4x = 25/10 = 2.5 4x = 10 ^ (2.5) x = 10 ^(2.5) / 4 = 79.056

OpenStudy (anonymous):

where does the log come in? What do you take the log of?

OpenStudy (anonymous):

jenn it was in your question

OpenStudy (anonymous):

First you solve for the log part. Then, to take the log off you use the relation: \[log_b(a) = k \iff b^k = a\]

OpenStudy (anonymous):

\[10\cdot log(4x) = 25\]\[\implies log(4x) = 2.5\]Since a 'bare' log is the log base 10..\[\implies log_{10}(4x) = 2.5\]\[\implies 10^{2.5} = 4x\]\[\implies 100\sqrt{10} = 4x\]\[\implies x = \frac{100}{4}\sqrt{10} = 25 \sqrt{10}\]

OpenStudy (anonymous):

I am confused. They are asking me for a numeric reply. Up to this point, we have not used the radical signs in this application. Nothing has been log AND radicals...can it be solved without the use of the radical?

OpenStudy (anonymous):

Sure, just raise 10 to the 2.5 power on your calculator.

OpenStudy (anonymous):

Then divide that answer by 4.

OpenStudy (anonymous):

Polpak is doing what someone has done in the beginning :-)

OpenStudy (anonymous):

OK - Thanks! I need the steps - I'm VERY VISUAL! I gave you BOTH medals - I appreciate the help!

OpenStudy (anonymous):

I'll start after I removed the log: \[\implies 10^{2.5} = 4x\]\[\implies 4x \approx316.228\]\[\implies x \approx \frac{316.228}{4}\approx79.057\]

OpenStudy (anonymous):

Can you help me with another one please? I will post it on the other board!

OpenStudy (anonymous):

Jenn.. you are most welcome... you can contact me in future for any help..

OpenStudy (anonymous):

Feel free to post more questions, but be sure you understand the principles of this problem first. Why did we do the things we did? How did it work? etc. Only then will you be able to do it yourself when it comes time to take the test.